OpenStudy (idealist10):

How to simplify e^[(-1)((2x+1)+lnabs(2x+1)+C)]?

OpenStudy (idealist10):

\[e ^{-1[(2x+1)+\ln \left| 2x+1 \right|+C]}\]

OpenStudy (dinamix):

\[K\frac{ 2x+1}{ e ^{2x+1}}\] @Idealist10

OpenStudy (dinamix):

that's only

OpenStudy (dinamix):

e^c = k

OpenStudy (dinamix):

@Idealist10 did u understand how i do it

OpenStudy (idealist10):

How did you get (2x+1)/(e^(2x+1))?

OpenStudy (dinamix):

\[e^{\ln(2x+1)}= 2x+1\] \[e ^{-(2x+1)}=\frac{ 1 }{ e ^{2x+1}}\]

OpenStudy (dinamix):

so ?

OpenStudy (idealist10):

But \[e ^{-\ln \left| 2x+1 \right|}\] is still 2x+1? Not 1/(2x+1)?

OpenStudy (dinamix):

no itd 1/2x+1

OpenStudy (freckles):

\[e^{- \ln|2x+1|}=e^{\ln|(2x+1)^{-1}|}=e^{\ln|\frac{1}{2x+1}|}=|\frac{1}{2x+1}|=\frac{1}{|2x+1|}\]